OpenStudy (anonymous):

solve: x/4=x-2/3. Please help mee

OpenStudy (nali):

x/4 = x- 2/3 4(x/4) = 4(x- 2/3) x = 4x - 8/3 -4x - 4x -3x = -8/ 3 (-1/3) -3x = (-1/3)-8/ 3 x = 8/9

OpenStudy (oaktree):

Subtract x/4 and add 2/3 to get 3/4x = 2/3. Solve to get x=8/9

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Wait that can't be right my answer choices are A. x = 4 B. x = 8 C. x = 2

OpenStudy (mertsj):

\[\frac{x}{4}\times12=12\times x-12\times \frac{2}{3}\] \[3x=12x-8\] \[-9x=-8\] \[x=\frac{8}{9}\]

OpenStudy (mertsj):

\[\frac{x}{4}=\frac{x-2}{3}\]Maybe we have misinterpreted the problem. Is it this:

OpenStudy (anonymous):

Yup

OpenStudy (mertsj):

So this is what you should have posted: x/4=(x-2)/3

OpenStudy (mertsj):

To solve it all you have to do is cross multiply

OpenStudy (anonymous):

Oh i see

OpenStudy (mertsj):

Did you get it?

OpenStudy (anonymous):

No